# ActSonic: Everyday Activity Recognition on Smart Glasses using Active Acoustic Sensing

### v1

@@ Abstract.
In this paper, we introduce ActSonic, an intelligent, low-power active acoustic sensing system integrated into eyeglasses. ActSonic is designed to recognize $27$ different everyday activities (e.g., eating, drinking, toothbrushing). It only needs a pair of miniature speakers and microphones mounted on each hinge of eyeglasses to emit ultrasonic waves to create an acoustic aura around the body. Based on the position and motion of various body parts, the acoustic signals are reflected with unique patterns captured by the microphone and analyzed by a customized self-supervised deep learning framework to infer the performed activities. ActSonic was deployed in a user study with 19 participants across 19 households to evaluate its efficacy. Without requiring any training data from a new user (leave-one-participant-out evaluation), ActSonic was able to detect 27 activities with an inference resolution of 1 second, achieving an average F1-score of 86.6% in an unconstrained setting and 93.4% in a prompted setting.
@@ 1. Introduction
Wearables are widely used globally to monitor daily behaviors and activities. Despite advancements in artificial intelligence, accurately tracking basic everyday human activities in real-world settings, often referred to as ”in the wild,” remains a significant challenge for these devices. For example, common wearables like smartwatches and glasses encounter difficulty in precisely tracking complex actions such as eating or drinking. The primary challenge in recognizing everyday human activities lies in consistently capturing high-quality information regarding different body parts involved in these activities.
Most wearable devices integrate inertial measurement units (IMUs), which, despite being small and low-power, often lack the necessary resolution for recognizing detailed human behaviors (Tong et al., 2020; Zhang et al., 2022). Researchers have explored wearable cameras (Price et al., 2022; Liu et al., 2023) as an alternative to improve activity identification, showcasing promising performance. However, integrating cameras into wearables faces practical limitations due to high energy consumption and significant data generation, raising privacy concerns in everyday contexts. Another approach utilizes microphones in wearables to capture sounds produced during various activities in the wild, offering low-power and cost-effective options. Yet, this method struggles with activities that do not generate distinct sounds (e.g., exercise, reading) (Iravantchi et al., 2021; Mollyn et al., 2022; Laput et al., 2018), constituting a significant portion of daily activities.
@@
Figure 1. Overview of the active acoustic sensing principle of ActSonic: The $x$-axis of the echo frames (in the 2nd row) represents the distance of echo reception. The corresponding video frames (in the first row) serve as activity references. The echo profile, created by stacking multiple echo frames, provides a spatio-temporal representation of the activity. These sliding windows with a duration of $2$ seconds of echo profiles (in the 3rd row) serve as inputs for the self-supervised learning algorithm.
Numerous activities in our daily lives involve the movements of different body parts, especially on the upper body and face. For instance, eating combines hand-to-mouth motion (using a spoon or chopsticks) with chewing. Therefore, activity recognition systems require detailed hand and mouth movement data to track fine-grained episodes of eating. Conversely, rinsing the mouth or yawning has very similar hand movements compared to eating. However, there are different movements of facial muscles if we follow the temporal progression of these activities. It is challenging to track the pose and movements of multiple body parts simultaneously at a single instrumentation point (e.g. IMU). As a result, existing wearable activity recognition systems suffer from a lack of precision and granularity in tracking complex activities in everyday life.
@@
In this paper, we introduce ActSonic, a self-supervised and low-power activity recognition system integrated into eyeglasses, based on active acoustic sensing. ActSonic is the first to demonstrate the feasibility of using active acoustic sensing on a wearable device to recognize 27 types of everyday activities without the need to collect any training data from a new user. Due to the low-power nature of acoustic sensors, it can operate for over 21 hours with a battery capacity equivalent to that of Google Glass (570 mAh). This work is motivated by the research question of creating a wearable sensing platform with a single instrumentation point to track a broad spectrum of everyday activities in real-world settings. It draws inspiration from recent advancements in using active acoustic sensing to monitor facial expressions (Li et al., 2022, 2024a) and upper body postures (Mahmud et al., 2023) via glasses or earphones, as these body postures are central to performing a variety of daily activities.
@@
We developed ActSonic by attaching a pair of miniature, low-power, off-the-shelf microphones and speakers to the hinges of glasses, respectively. The sensing system emits inaudible ultrasonic waves to create an acoustic aura around the body. Based on the shape and position of various body parts, the acoustic signals are reflected with unique patterns captured by the microphone. We developed a customized self-supervised deep learning framework to interpret the reflected signals, which are presented with complex multipath echoes and include rich information about movements on both the face and upper body, to infer the performed activities.
@@
ActSonic was evaluated comprehensively in two studies in real-world settings. The first study was a semi-in-the-wild investigation involving $12$ participants. In this study, each participant performed all 27 activities at their homes in the presence of a researcher. To further validate the system’s performance in completely uncontrolled real-world conditions, we conducted a second study with $7$ participants. In this study, participants were provided with the device at their homes to record their unconstrained daily activities alone without any intervention. These two studies resulted in the collection of $40$ hours of activity data from $19$ different households. The leave-one-participant-out evaluation showed that ActSonic achieved an average F1-score of $93.4\%$ in the first semi-in-the-wild user study and $86.6\%$ in the second in-the-wild study.
ActSonic has significantly advanced wearable-based activity recognition, demonstrating promising performance across various dimensions. Unlike previous data-driven systems that require user-specific training data, ActSonic accurately recognizes 27 activities in participants’ homes without the need for individualized training data collection. Additionally, it offers an affordable and low-power hardware solution, enabling over 21 hours of continuous operation on wearables with small batteries—a notable departure from high-power signature sensing systems, such as cameras. While few wearable-based activity recognition systems have been evaluated in uncontrolled scenarios, ActSonic demonstrates robust performance in recognizing a diverse set of 27 activities in real-world settings. Moreover, compared to many previous works relying on multiple sensors for limited activities, ActSonic achieves the recognition of 27 activities with a single device, significantly reducing barriers to tracking everyday activities using glasses in real-world contexts.
In summary, the contributions of the paper are:
@@
  • The first demonstration of utilizing low-power active acoustic sensing for fine-grained activity recognition.
  • We developed a self-supervised deep learning framework that is able to distinguish $27$ activities from the received reflected acoustic signals
  • We conducted a semi-in-the-wild study and an in-the-wild study with $19$ participants at $19$ homes to collect over $40$ hours of activity data, demonstrating a promising performance with F1 scores of $93.4\%$ and $86.6\%$ respectively
@@ 2.3. Acoustic Sensing-based Human Activity Recognition
The aforementioned vision-based and multimodal sensing approaches for detecting human activities offer higher spatiotemporal resolution, leading to lower recognition errors. However, the usage of multiple sensors raises concerns about increased power consumption and privacy invasion. Recently, researchers have leveraged the pervasiveness of microphones in off-the-shelf commodity devices to recognize human activities (Wu et al., 2020; Liang and Thomaz, 2019; Laput et al., 2018; Lane et al., 2015). These acoustic sensing systems utilize passively sensed audio within the audible frequency range ($20$ Hz to $16$ KHz). Although passively sensed environmental sound provides discriminative information required to infer certain activities that generate sound, it raises serious privacy concerns since it may record personal conversations. In response to this, recent works (Mollyn et al., 2022; Chen et al., 2008) adopt subsampling and other preprocessing techniques to make the audio unintelligible and then recognize activities based on that. Additionally, passively sensed audio from inaudible frequency ranges (infrasonic and ultrasonic) has been utilized to recognize daily activities (Kijima et al., 2018; Iravantchi et al., 2021, 2019). While these techniques offer better preservation of user privacy, activity recognition on these systems relies on the assumption that activities will generate environmental sound that can be modeled by the system. However, many activities in daily living involve body-limb movement and do not necessarily generate distinctive sounds. Active acoustic sensing-based approaches emit high-frequency sound waves and leverage the reflected sounds to capture fine-grained facial movements (Li et al., 2022; Zhang et al., 2023; Sun et al., 2023; Li et al., 2024b), hand poses (Lee et al., 2024), sign language gesture (Jin et al., 2021, 2023), body pose (Mahmud et al., 2023), gaze (Li et al., 2024c)and vital signs (Wang et al., 2018; Nandakumar et al., 2015).
@@ 3. Design and Implementation of Sensing System
Our goal is to develop a wearable activity recognition platform integrated into glasses, enabling the identification of a wide range of everyday activities based on the positions and movements of different body parts, as captured by active acoustic sensing technology. While prior research (Mahmud et al., 2023; Li et al., 2022; Zhang et al., 2023; Li et al., 2024a) has underscored the efficacy of active acoustic sensing in detecting facial changes and upper body poses, no existing system has seamlessly integrated this sensing modality for the comprehensive task of classifying diverse daily activities in real-world scenarios. Given that most daily activities entail movements across various body parts, particularly the upper body, head, and face, our primary focus with ActSonic is to capture and differentiate these activities based on the position and movement of these body areas, as detected by our sensing system. ActSonic accomplishes this with a singular device, facilitating concurrent information capture. This section provides an overview of the active acoustic sensing setup, the process of feature extraction for activity recognition, and the hardware implementation, with a particular emphasis on the wearable form factor of the system.
@@ 3.1. Configuration of Active Acoustic Signal
The active acoustic sensing system in ActSonic integrates two pairs of ultrasonic transmitters and receivers on eyeglass hinges. Utilizing Frequency Modulated Continuous Wave (FMCW) chirps, ActSonic emits ultrasonic signals with linearly modulated frequencies ranging from $18$ to $21.5$ KHz and $21.5$ to $24.5$ KHz for the left and right transmitters, respectively, each with a bandwidth of $B=3$ KHz. The ActSonic system samples these FMCW chirps at $f_{s}=50$ KHz. By employing cross-correlation-based FMCW (Wang et al., 2018), the system achieves a minimum discernible distance, or resolution, of $\frac{c}{2{f_{s}}}=0.343$ cm ($c=343$ m/s in dry air at $20\degree C$). With a sweep period $T$ of $12$ ms for transmitted FMCW chirps (comprising $N=600$ samples), ActSonic’s maximum sensing range extends to approximately $2$ meters. This combination of high sensing resolution ($0.343$ cm) and extensive sensing range enables us to capture both subtle skin deformations on the face and monitor the pose and coarse movement of the upper body region effectively.
@@ 3.2. Computation of Echo Profile and Acoustic Flow
The cross-correlation matches the sweep period of the emitted FMCW ultrasonic wave. It generates an echo frame, represented as a $(600\times 1)$ column vector in ActSonic. Stacking these frames creates the echo profile (Nandakumar et al., 2016; Wang et al., 2018; Li et al., 2022; Zhang et al., 2023), where brighter pixels indicate strong reflections at specific distances. With two transmitter-receiver pairs, ActSonic accounts for four transmission paths. To elaborate, if we consider the left and right transmitter-receiver pair as $(T_{left},R_{left})$ and $(T_{right},R_{right})$ respectively, then the paths will be $T_{left}\rightarrow R_{left}$, $T_{left}\rightarrow R_{right}$, $T_{right}\rightarrow R_{left}$, and $T_{right}\rightarrow R_{right}$. In ActSonic, we stack the outputs of cross-correlation of these four paths as four channels of the echo profile. Note that the computation of one channel in the echo profile is shown in Figure 2.

### v2

# ActSonic: Recognizing Everyday Activities from Inaudible Acoustic Waves Around the Body

@@ Abstract.
We present ActSonic, an intelligent, low-power active acoustic sensing system integrated into eyeglasses that can recognize $27$ different everyday activities (e.g., eating, drinking, toothbrushing) from inaudible acoustic waves around the body with a time resolution of one second. It only needs a pair of miniature speakers and microphones mounted on each hinge of eyeglasses to emit ultrasonic waves to create an acoustic aura around the body. Based on the position and motion of various body parts, the acoustic signals are reflected with unique patterns captured by the microphone and analyzed by a customized self-supervised deep learning framework to infer the performed activities. ActSonic was deployed in a user study with 19 participants across 19 households to evaluate its efficacy. Without requiring any training data from a new user (leave-one-participant-out evaluation), ActSonic was able to detect 27 activities, achieving an average F1-score of 86.6% in fully unconstrained scenarios and 93.4% in prompted settings at participants’ homes.
@@ 1. Introduction
Figure 1. Overview of the active acoustic sensing principle of ActSonic: The $x$-axis of the echo frames (in the 2nd row) represents the distance of echo reception. The corresponding video frames (in the first row) serve as activity references. The echo profile, created by stacking multiple echo frames, provides a spatiotemporal representation of the activity. These sliding windows with a duration of $2$ seconds of echo profiles (in the 3rd row) serve as inputs for the self-supervised learning algorithm.
Despite smart glasses becoming popular and being worn daily, existing eyeglasses like many other commodity wearables, still lack the ability to understand the user’s fine-grained everyday activities in real-world settings. Distinguishing human activities is challenging due to the substantial variance across different activities, each involving complex and subtle movements of multiple body parts (e.g., face, arms, hands). For example, eating requires the coordination of fingers and arms to deliver food to the mouth, as well as the movement of facial muscles for chewing.
Acquiring precise body pose data from various body parts is essential for accurately distinguishing different human activities. However, achieving this with everyday wearable devices (e.g., eyeglasses) has presented persistent challenges: Wearable cameras (Price et al., 2022; Liu et al., 2023) consume significant power, rapidly draining batteries. Furthermore, transmitting and processing video data on wearable devices is resource-intensive. Many sensing systems (Inertial Measurement Units (IMUs) (Tong et al., 2020; Zhang et al., 2022) or Electromyography (EMG) (Zhang and Amft, 2017, 2018) ) only capture information around the instrumentation position, which is insufficient for activities involving multiple body parts (e.g., face and hands) unless multiple devices are worn on the body, which is less preferred by many users. Microphone-recorded passive sounds (Iravantchi et al., 2021; Mollyn et al., 2022; Laput et al., 2018) have shown promising performance in tracking human activities in the wild. However, it does not work effectively for activities lacking distinctive auditory cues. Therefore, achieving accurate recognition of each nuanced activity often necessitates a new set of customization on wearables with carefully selected sensor combinations, a solution that may not generalize well across different activities. Unlike the computer vision community, which primarily relies on cameras as the sole type of sensors to record data, facilitating the establishment of various types of benchmark datasets, researchers in the wearable and Ubicomp community face challenges in collecting data to track fine-grained human behaviors.
In this paper, we aim to answer the question: Is it possible to develop a simple sensing system on eyeglasses that is low-cost, minimally-obtrusive, energy-efficient, and privacy-aware, while capable of recognizing a wide range of everyday human activities? A positive answer would immediately lower the barriers for researchers to study and track human activities of their interest.
To tackle this challenge, we utilized an empirical observation of human activities that a significant portion of human activities involve both facial movements and upper body movements. If a single sensing system on eyeglasses can simultaneously capture high-quality facial movements and upper body poses, this captured information can potentially be learned by a customized deep learning framework to distinguish a wide set of human activities. However, designing such a wearable system that can accurately recognize various types of human activities with high resolution is highly challenging:
  • The system must operate on low power to sustain continuous usage throughout the day.
  • Privacy concerns of the user on wearing such a wearable system daily can be potentially alleviated with a carefully designed solution.
  • The appearance of the wearable needs to be largely unchanged while wearing comfort is not compromised.
  • The system must demonstrate robustness to accommodate various daily activities and distinguish them from other categories, including ”Null”, which refers to activities not included in the tracking set.
  • Demonstrating the efficacy of the system with high resolution requires recording and labeling ground truth data at the second level, a challenging and time-consuming task that many prior works overlooked.
Recent research has demonstrated the feasibility of integrating active acoustic sensing into eyeglasses using commodity microphones and speakers. By analyzing the reflected acoustic waves from the eyes, face, or upper body, it becomes possible to infer gaze (Li et al., 2024c), facial expression (Zhang et al., 2023b; Li et al., 2024a), or upper body pose (Mahmud et al., 2023; Jin et al., 2021) respectively. This development underscores the potential of leveraging a single sensing modality—active acoustic sensing on eyeglasses—to simultaneously capture both facial movements and upper body poses, which can be used for activity recognition. In this paper, we aim to answer the research question:
  • Can we utilize the inaudible active acoustic signals reflected around the human body to recognize a rich set of everyday human activities?
Given the small size, relatively low power consumption, low cost, and ubiquity of microphones and speakers, such eyeglasses capable of tracking diverse human activities would significantly reduce the barriers to activity monitoring, a critical task within the Ubicomp community. Instead of relying on multiple devices for different activities, this all-in-one system could effectively track and differentiate between multiple activities simultaneously. However, designing such an active-acoustic sensing system and proving its efficacy for everyday activity recognition presents significant challenges:
  • While active acoustic sensing has demonstrated effectiveness in recognizing facial expressions and upper body pose respectively, it remains unclear if the fused reflected acoustic signals provide adequate information to distinguish a rich set of fine-grained human activities.
  • Designing machine learning and signal processing algorithms that focus on temporal and structural differences in activities from these reflected acoustic signals, rather than individual body shapes, for usability to minimize user effort for training.
  • Robustness to various acoustic noises, such as multi-path acoustic reflections from the environmental setups (e.g., surrounding furniture) and potential overlap with the ultrasound band, is essential for handling different environmental conditions.
  • Capturing nuanced variations in activities, particularly in large motions where finer details are significant, is crucial. For example, recognizing activities with similar arm or facial movements (e.g., yawning vs. coughing) requires an understanding of the combination of movements and temporal patterns.
  • Robustness towards a large number of unknown or unlabeled activities in the ”Null” category is necessary.
@@
To address this research question, we developed ActSonic, a self-supervised and low-power activity recognition system integrated into eyeglasses, utilizing active acoustic sensing. It is the first system to demonstrate the feasibility of recognizing 27 types of everyday activities without collecting training data from new users. Thanks to the low-power nature of acoustic sensors, ActSonic can operate for over 21 hours with a battery capacity equivalent to that of Google Glass (570 mAh).
@@
We developed ActSonic by attaching a pair of miniature, low-power, off-the-shelf microphones and speakers to the hinges of glasses, respectively. The sensing system emits inaudible ultrasonic waves to create an acoustic aura around the body. Based on the shape and position of various body parts, the acoustic signals are reflected with unique and fused patterns captured by the microphone. We developed a customized self-supervised deep learning framework to interpret the reflected signals at areas of interest and their temporal patterns, which are presented with complex multipath echoes and include rich information about movements on both the face and upper body, to infer the performed activities.
@@
ActSonic was evaluated comprehensively in two studies in real-world settings. The first study was a semi-in-the-wild investigation involving 12 participants. In this study, each participant performed all 27 activities at their own homes (12 homes) in the presence of a researcher. To further validate the system’s performance in completely uncontrolled real-world conditions, we conducted a second study with 7 participants. In this second study, participants were provided with the device at their homes to record their unconstrained daily activities alone without any intervention. These two studies resulted in the collection of 40 hours of activity data from 19 different households, which were labeled with ground truth at each second. The leave-one-participant-out evaluation showed that ActSonic was able to distinguish these 27 activities at these 19 homes at each second without any training data from the user, with an average F1-score of 93.4% in the first semi-in-the-wild user study and 86.6% in the second in-the-wild study.
ActSonic advance the field of eyeglasses-based activity recognition and lower the barrier of tracking everyday human activities in the following aspects:
@@
  • The first demonstration of creating and utilizing inaudible acoustic waves around the body for fine-grained everyday activity recognition.
  • Unlike many new data-driven activity recognition systems on wearables, ActSonic achieved promising performance without requiring training data from a new user or a new environment, which makes collecting human activity data much easier.
  • Microphones and speakers are readily available, cost-effective to set up, and can sustain prolonged operation on a small battery attached to glasses, facilitating easy replication of our system by other researchers.
  • Compared to many previous works relying on multiple sensors or instrumentations, ActSonic accurately recognizes 27 activities with a single device at a high time resolution of one second.
@@ 2.3. Acoustic Sensing-based Human Activity Recognition
The aforementioned vision-based and multimodal sensing approaches for detecting human activities offer higher spatiotemporal resolution, leading to lower recognition errors. However, the usage of multiple sensors raises concerns about increased power consumption and privacy invasion. Recently, researchers have leveraged the pervasiveness of microphones in off-the-shelf commodity devices to recognize human activities (Wu et al., 2020; Liang and Thomaz, 2019; Laput et al., 2018; Lane et al., 2015). These acoustic sensing systems utilize passively sensed audio within the audible frequency range ($20$ Hz to $16$ KHz). Although passively sensed environmental sound provides discriminative information required to infer certain activities that generate sound, it raises serious privacy concerns since it may record personal conversations. In response to this, recent works (Mollyn et al., 2022; Chen et al., 2008) adopt subsampling and other preprocessing techniques to make the audio unintelligible and then recognize activities based on that. Additionally, passively sensed audio from inaudible frequency ranges (infrasonic and ultrasonic) has been utilized to recognize daily activities (Kijima et al., 2018; Iravantchi et al., 2021, 2019). While these techniques offer better preservation of user privacy, activity recognition on these systems relies on the assumption that activities will generate environmental sound that can be modeled by the system. However, many activities in daily living involve body-limb movement and do not necessarily generate distinctive sounds. Active acoustic sensing-based approaches emit high-frequency sound waves and leverage the reflected sounds to capture fine-grained facial movements (Li et al., 2022; Zhang et al., 2023a; Sun et al., 2023; Li et al., 2024b), hand poses (Lee et al., 2024), sign language gesture (Jin et al., 2021, 2023), body pose (Mahmud et al., 2023), gaze (Li et al., 2024c)and vital signs (Wang et al., 2018; Nandakumar et al., 2015).
@@ 3. Design and Implementation of Sensing System
Our goal is to develop a wearable activity recognition platform integrated into glasses, enabling the identification of a wide range of everyday activities based on the positions and movements of different body parts, as captured by active acoustic sensing technology. While prior research (Mahmud et al., 2023; Li et al., 2022; Zhang et al., 2023a; Li et al., 2024a) has underscored the efficacy of active acoustic sensing in detecting facial changes and upper body poses, no existing system has seamlessly integrated this sensing modality for the comprehensive task of classifying diverse daily activities in real-world scenarios. Given that most daily activities entail movements across various body parts, particularly the upper body, head, and face, our primary focus with ActSonic is to capture and differentiate these activities based on the position and movement of these body areas, as detected by our sensing system. ActSonic accomplishes this with a singular device, facilitating concurrent information capture. This section provides an overview of the active acoustic sensing setup, the process of feature extraction for activity recognition, and the hardware implementation, with a particular emphasis on the wearable form factor of the system.
@@ 3.1. Configuration of Active Acoustic Signal
The active acoustic sensing system in ActSonic integrates two pairs of ultrasonic transmitters and receivers on eyeglass hinges. Utilizing Cross-correlation based Frequency Modulated Continuous Wave (C-FMCW) (Wang et al., 2018) chirps, ActSonic emits ultrasonic signals with linearly modulated frequencies ranging from $18$ to $21.5$ kHz and $21.5$ to $24.5$ kHz for the left and right transmitters, respectively, each with a bandwidth of $B=3$ kHz.
For the C-FMCW (Wang et al., 2018) technique employed by ActSonic, the theoretical range resolution $R$ can be derived as follows:
(1) $R=\frac{C\cdot Lag}{2F_{s}}-vt$
where $C$ is the speed of sound ($C=343$ m/s in dry air at $20\degree$C), $F_{s}$ is the sampling frequency of the microphone or receiver ($F_{s}=50$ kHz in the case of ActSonic), $v$ is the velocity of the tracked body part, $t$ is the elapsed time from the start of the modulation period, and $Lag$ is the number of samples shifted between the transmitted and received acoustic signals. With $N=600$ samples transmitted for each chirp, ActSonic’s sensing system has a sweep period of $T=\frac{600}{50000}=0.012$ seconds $=12$ milliseconds. Thus, the value of $t$ ranges from $0\leq t\leq 0.012$ s, and the value of $Lag$ in Eq. 1 ranges from $0\leq Lag\leq 600$.
According to the rationale provided in (Wang et al., 2018), we can set $v=0$ in Eq. 1 in the case where the tracked body part is static or moving slowly with respect to the speed of sound. The velocity of extremely fast arm movements, such as punching, has been measured at approximately $8.9$ to $11.5$ m/s (Kimm and Thiel, 2015), while the duration of the fastest facial expression in humans is measured at 100 ms (Yan et al., 2013). The speed of these extreme body motions ($v$) is significantly smaller compared to the speed of sound ($C$). Therefore, based on the equation above, the impact of the sensing resolution ($vt$) caused by body movements is minimal. As a result, the theoretical range resolution $\delta R$ of the C-FMCW signal deployed in ActSonic can be simplified to:
(2) $\delta R=\frac{C\cdot\delta Lag}{2F_{s}}$
Since cross-correlation is computed for each sample in the case of C-FMCW (Wang et al., 2018), we can set $\delta Lag=1$ for the theoretical range resolution calculation in Eq. 2. Therefore, we can substitute the values of variables in Eq. 2 and compute the resolution of the sensing system of ActSonic as follows:
(3) $\delta R=\frac{343\text{ m/s}}{2\cdot 50000}=0.343\text{ cm}=3.43\text{ mm}$
The maximum sensing range of ActSonic can be computed as $R_{\text{max}}=N\cdot\delta R=600\cdot 0.343\text{ cm}=205.8\text{ cm}\approx 2\text{ m}$. This combination of high sensing resolution ($0.343$ cm) and extensive sensing range enables us to capture both subtle skin deformations on the face and monitor the pose and coarse movement of the upper body region effectively. Additionally, it’s important to note that the sensing resolution mentioned above is calculated for range-finding applications, as demonstrated in (Wang et al., 2018), which differs from our approach. This calculation is provided as a reference to facilitate readers’ understanding of the sensing principle underlying our proposed system.
@@ 3.2. Computation of Echo Profile and Acoustic Flow
The cross-correlation matches the sweep period of the emitted FMCW ultrasonic wave. It generates an echo frame, represented as a $(600\times 1)$ column vector in ActSonic. Stacking these frames creates the echo profile (Nandakumar et al., 2016; Wang et al., 2018; Li et al., 2022; Zhang et al., 2023a), where brighter pixels indicate strong reflections at specific distances. With two transmitter-receiver pairs, ActSonic accounts for four transmission paths. To elaborate, if we consider the left and right transmitter-receiver pair as $(T_{left},R_{left})$ and $(T_{right},R_{right})$ respectively, then the paths will be $T_{left}\rightarrow R_{left}$, $T_{left}\rightarrow R_{right}$, $T_{right}\rightarrow R_{left}$, and $T_{right}\rightarrow R_{right}$. In ActSonic, we stack the outputs of cross-correlation of these four paths as four channels of the echo profile. Note that the computation of one channel in the echo profile is shown in Figure 2.